### v2

# Anti-Phishing Training (Still) Does Not Work: A Large-Scale Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale

Abstract
Social engineering attacks delivered via email, commonly known as phishing, represent a persistent cybersecurity threat leading to significant organizational incidents and data breaches. Although many organizations train employees on phishing, often mandated by compliance requirements, the real-world effectiveness of this training remains debated.
To contribute to evidence-based cybersecurity policy, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm. Our experimental design refined prior work by comparing training modalities in operational environments, validating NIST’s standardized phishing difficulty measurement, and introducing novel organizational-level temporal resilience metrics. Echoing prior work, training interventions showed no significant main effects on click rates (pp=0.450) or reporting rates (pp=0.417), with negligible effect sizes. However, we found that the NIST Phish Scale predicted user behavior, with click rates increasing from 7.0% for easy lures to 15.0% for hard lures. Our organizational-level resilience result was mixed: 36-55% of campaigns achieved “innoculation” patterns where reports preceded clicks, but training did not significantly improve organizational-level temporal protection. In summary, our results confirm the ineffectiveness of current phishing training approaches while offering a refined study design for future work.
1 Introduction
Phishing attacks harm organizations through operational incidents, data breaches, and financial losses [4]. Industry studies and the US FBI agree that phishing is the primary attack vector to breach organizations [42]. As an example, in 2022–2023 campaigns attributed to the Scattered Spider cluster (aka 0ktapus/Muddled Libra) compromised 9,931 accounts in over 130 organizations [43, 23]. This campaign cost MGM Resorts alone an estimated \$100 million [79]. To prevent this class of attacks, organizations employ technical controls (e.g., spam filters, attachment analyzers) complemented by cybersecurity awareness training [58, 48].
Despite widespread organizational investment in anti-phishing training, mounting empirical evidence (and employee sentiment [78]) questions its effectiveness. Large-scale studies across multiple sectors demonstrate minimal training impact [48, 53, 54]. Most pertinently, Lain et al. showed that phishing training did not improve security outcomes across 19,000+ employees over multiple years [53, 54]. These findings align with theoretical work by Volkamer et al., who demonstrate fundamental limitations in human processing of security indicators [96].
This growing body of evidence suggests that phishing training is ineffective, but we believe that further reproduction of the result is needed to inform cybersecurity policy and defense strategies [67, 89]. [^1] Our reproduction effort differs from an exact replication, which would involve repeating identical procedures. Instead we focus on testing the generalizability of the phenomenon of training ineffectiveness, varying the organizational context and the training approach. We target two specific knowledge gaps. First, existing studies lack standardized measures of phishing difficulty, so we do not know whether training effectiveness varies by attack sophistication. Second, most studies examine single training modalities rather than comparing different approaches systematically, so we do not know if some training designs are more effective than others.
To address these limitations, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm. Our study design has three distinctive elements: (1) mandatory pre-phishing training using two distinct training modalities (lecture-based vs. interactive); (2) systematic difficulty assessment using the NIST Phish Scale [86, 7]; and (3) novel organizational temporal resilience measurement through the Organizational Inoculation Index, which measures the effectiveness of phishing training from an organizational perspective rather than an individual one. This design enables rigorous evaluation of training effectiveness across standardized difficulty levels while providing new insights into organizational cybersecurity resilience patterns that complement individual-focused assessments.
Our reproduction study confirms the ineffectiveness of anti-phishing training across multiple dimensions, but adds nuance to prior results. Consistent with prior work, neither training modality produced statistically significant improvements in phish click reduction (pp=0.450) or phish reporting behavior (pp=0.417), with negligible effect sizes indicating minimal practical value for organizational security investment. However, the NIST Phish Scale successfully predicted user behavior (F(2, 12086) = 41.415, p < 0.001, $\eta^{2}$ = 0.007), providing the first large-scale validation of this standardized difficulty measure. [^2] While the effect size is small in statistical terms due to low base rates of phishing engagement, the practical implications are significant: click rates doubled from 7.0% to 15.0% as lure difficulty increased. Beyond that, our Organizational Inoculation Index revealed substantial organizational resilience that operates independently of individual training effectiveness. Across all experimental conditions, 36.4% to 54.5% of template-group combinations demonstrated “inoculation” patterns where threat reports (from a security-conscious employee) preceded threat interaction by another employee.
We summarize our main contributions below:
  1. 1.
    Reproduction of Training Ineffectiveness Across Multiple Modalities. Our results reproduce the core finding of training ineffectiveness while demonstrating this phenomenon’s generalizability across training designs.
  2. 2.
    First Large-Scale Validation of the NIST Phish Scale. We evaluate the NIST Phish Scale across 12,511 participants. The Phish Scale is a viable measure for assessing phishing difficulty, with predictive utility.
  3. 3.
    Introduction of the Organizational Inoculation Index. We introduce and validate a novel metric for organizational temporal security. We report substantial organizational resilience (36-55% inoculation rates), which operates independent of individual training effectiveness and provides new insights into collective cybersecurity behaviors.
Significance: Cybersecurity practice and policy must be evidence-based. For regulators, our work contributes to growing evidence that compliance-driven training requirements may not achieve intended security outcomes. For organizations, our results suggest realistic (low) expectations about training outcomes and highlight the importance of technical controls rather than human-centered defenses. However, we also demonstrated that collective security behaviors create measurable protective effects, emphasizing the value of feedback loops between user reports and real-time threat response. For researchers, we validate NIST’s standardized difficulty measure and confirm the reproducibility of training ineffectiveness findings across organizational contexts.
2 Background and Related Work
Figure 1: Phishing defense architecture, mapping the email pathway from authentication through post-delivery response. Technical controls create multiple interception points prior to human interaction, and human reports provide a feedback opportunity for organizational-level threat mitigation. See Appendix 2 for the full control pathway in the studied organization.
We review phishing attacks and defenses (§2.1), and then discuss prior empirical studies on phishing training (§2.2).
2.1 Phishing Attacks and Defenses
Phishing attacks pose substantial risks to organizations, who erect both technical and human defenses against them.
2.1.1 Phishing Attacks
In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests, and the interests of their organizations. Attackers seek results such as sensitive information, account credentials, or the ability to control the victim’s machine [4, 34]. Traditionally, phishing attacks are divided into two classes [25]:
Mass Phishing Campaigns: These use malicious emails sent to thousands or millions of recipients, with common themes such as banking, shipping, and software updates [4]. Individual success rates are low (1-3%) [3], but the massive scale makes them profitable.
Spear Phishing and Targeted Attacks: Spear phishing targets specific individuals or organizations with personalized content from open-source intelligence (OSINT) and reconnaissance. Attackers mine sources such as social media to craft messages that reference real relationships and events [13]. These tailored campaigns have higher success rates (10–20%) [14]. Successful spear phishing attacks include the 2016 Democratic National Committee breach and widespread business email compromise (BEC) schemes [4].
Modern Phishing: Contemporary phishing operations leverage infrastructure including compromised (“trusted”) websites and domain generation algorithms to maintain persistence and evade takedown efforts [3]. In addition, Generative AI is enabling both mass phishing and spear phishing through targeted and fluent emails and other media [4, 80]. LLM-generated phishing content now achieves effectiveness comparable to human experts in organizational settings [9].
2.1.2 Automated Defenses
Figure˜1 illustrates how these attacks flow through an organization’s email system. Many technical defenses (blue boxes) combat attackers’ behavior, such as SPF, DKIM, and DMARC to verify that mail comes unaltered from authorized infrastructure [51, 22, 52]; ARC for preserving authentication results across forwarding [5]; BIMI for verified logos [10, 12]; and secure email gateways (SEG) for content and header inspection and URL rewriting for time-of-click inspection [1].
After delivery, organizations employ continuous monitoring and feedback loops to adapt controls to evolving campaigns. Organizations blend threat-intelligence updates with AI/ML-driven anomaly scoring and impersonation detection to surface novel lures [58]. For example, security operations center (SOC) analysts triage user-reported messages, promote high-confidence indicators back into blocklists, and tune SEG policies based on observed false positives/negatives. Because adversaries adapt, detection is a dynamic process [4, 3].
See Appendix 2 for the details of the studied organization’s email controls and how user phishing reports create feedback loops that strengthen automated defenses.
2.1.3 Phishing Training as a Defense
Because phishing attacks compromise these technical defenses [14, 13], the human recipients play an important role in phishing defense [30, 15, 2, 28]. There are many approaches to training humans about phishing attacks. Methods range from lecture-style training involving passive videos and readings [56] to simulation feedback providing immediate feedback after interaction with phishing simulations [19, 18]. Interactive training uses gamification or scenario-based learning to engage participants [11, 46], while just-in-time training alerts users in real time during risky behavior [36].
Anti-phishing trainings are often included in the cybersecurity awareness trainings ubiquitous in organizations. Legal mandates drive this practice, e.g., HIPAA requiring security awareness training for healthcare organizations [74]. GDPR and PCI DSS encourage or mandate similar training [76, 21]. Industry standards and frameworks further reinforce these requirements [75], e.g., ISO 27001 calling for human resource security controls including awareness programs [35].
2.2 Empirical Studies of Phishing Training
Table 1: Summary of Prior Experiments on Phishing and Training Interventions. Studies are organized by research context (Lab vs. Real-World). Context: “Lab” denotes research simulations, where subjects are not in their work roles. “Real-World” denotes studies where participants were employees acting in their job capacity. Complexity measure: Whether the study controlled for the difficulty of the phishing tasks. indicates the use of an open control such as NIST Phish Scale, and indicates no control. Hypotheses: How these works informed the hypotheses in the present study (cf. Table˜4).
Work Context Population Sample size Complexity Measure Compared Training Approaches? Hypotheses
[82] (2024) Lab University students 96 H3
[16] (2024) Lab University students 117 (Phish Scale) H4
[18] (2014) Real-world Employees 1,359 H2, H4
[19] (2017) Real-world University students 150 H2, H3
[90] (2022) Real-world Employees 119 H2
[53] (2022) Real-world Employees 19,000+ H2
[47] (2023) Real-world Finance employees 5,000 H2
[54] (2024) Real-world Employees 4,554 H2, H3
[29] (2024) Real-world Banking employees “Thousands” H2
[48] (2025) Real-world Healthcare employees 19,500+ H2
Ours Real-world Fintech org. 12,511 (Phish Scale)
Our review and analysis of prior work is summarized in Table˜1. We contribute to the body of knowledge on phishing training by conducting another large-scale study situated in a real-world context (table columns 2-4), incorporating a complexity measure (column 5), and a comparison of multiple training approaches (column 6). Our study tests hypotheses derived from prior work (final column). The remainder of this section describes these contributions. For general surveys of knowledge about human factors in phishing training, we refer the reader to Marshall et al. [56] and Franz et al. [37].
Lab vs. Real-world Studies: Many early phishing training studies were conducted in controlled laboratory settings, where researchers benefit from a high degree of experimental control. Studies of this sort have provided numerous insights, including that active learning and immediate feedback can improve short-term detection rates, and have explored how individual differences affect training outcomes [90, 56]. However, as noted by Franz et al. [37], these controlled settings may lack ecological validity. In laboratory environments, participant samples are typically small and consist of homogeneous groups such as academics and students.
In response, a growing number of studies have begun examining phishing training in real-world organizational environments. For example, Doing et al. [29] analyzed phishing behaviors in bank employees, and Ho et al. [48] studied healthcare employees. Lain et al. has also conducted large-scale real-world measurements of training effectiveness [53, 54] Our study echoes some of these prior results in a different organizational context.
Controlling for Lure Complexity: These studies consistently report little effectiveness of anti-phishing training. These works have not, however, incorporated a standardized method for assessing the difficulty of the phishing lures they use [48, 29]. We contribute the first real-world measurement of the effect of lure complexity on subjects’ interaction behaviors, based in NIST’s Phish Scale.
To assess phishing lure complexity, the NIST Phish Scale categorizes them along two dimensions (Table˜2). Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities [16]. Premise Alignment measures the relevance and familiarity of the email’s subject matter to the recipient’s typical email environment and organizational context [6]. Higher alignment messages with fewer cues are the most deceptive [14]. The premise of Phish Scale is to allow organizations to benchmark lure complexity and contextualize simulation outcomes [17]. Barrientos et al. [8] conducted early validation work on the NIST Phish Scale, while Dawkins and Jacobs [27] provided the official implementation guidelines. Ours is the first study to validate the Phish Scale’s premise at scale.
Training Approaches: Franz et al. [37] taxonomize five categories of phishing interventions: training, awareness, education, notifications, and design interventions. Most studies have focused on training, which can be divided into lecture-style annual training and embedded training (e.g., through ongoing phishing simulations sent by the organization’s security team). Like some prior studies, our study performs a controlled experiment to compare the effectiveness of two forms of phishing training.
Table 2: NIST Phish Scale for Detection Difficulty [87, 27].
Cues \ Alignment Weak Medium Strong
Few Medium Hard Hard
Some Easy-to-Medium Medium Hard
Many Easy Medium Medium
This study addresses those gaps by applying the NIST Phish Scale [27] to evaluate training effectiveness across varied phishing difficulty levels in an enterprise deployment with over 12,000 participants. In this study, our objectives are to: (1) operationalize the NIST framework at scale, (2) examine how difficulty influences detection and reporting across training modalities, (3) provide comparative evidence on lecture-based versus interactive approaches, and (4) offer guidance for large-scale awareness efforts under conditions of modest effectiveness [48]. This enables evidence-based decision-making on phishing defense strategies that account for both training design and message complexity.
3 Methodology
To summarize the preceding analysis: phishing training is widely practiced, though recent evidence sheds doubt on its utility. Few studies have rigorously assessed outcomes against standardized difficulty measures at sufficient scale for robust inference. We provide a reproduction of this observation, tested through hypotheses derived from prior works.
In this section, we describe our methodology: study design, subjects, hypotheses, metrics, and analysis methods. For ethics considerations, see the associated appendix.
3.1 Study Design
Figure 2: Illustration of our methodology. Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects. We compared between-subject performance using a mix of standard and novel metrics to test our five hypotheses.
Our study design is illustrated in Figure˜2. We employed a two-factor between-subjects experimental design [47]. Subjects were randomly assigned to one of three experimental conditions using our organization’s training management system. Each condition reflected the phishing training they received. Subjects were randomly selected, trained, and later received simulated phishing emails of varying complexity. We quantitatively compared subjects’ performance in handling these phishing emails between the treatments.
Factor 1—Training: The research team collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment. This partnership enabled access to professional training content and established phishing templates.
The vendor offers two distinct training approaches for phishing. The first approach, Treatment A, consists of lecture-based videos on the dangers of various social engineering attacks, including phishing recognition tactics, social engineering techniques, and organizational security protocols. Figure˜3 gives a screenshot. Comprehension is assessed through a quiz. The second approach, Treatment B, consists of those videos followed by a series of interactive phishing exercises that present users with samples from a library of approximately 20 phishing templates. These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale, and reinforces appropriate response behaviors such as the “Report Phish” button in the email client [19]. Finally, we included a Control Group whose members received no cybersecurity training prior to the phishing simulation phase. This group was included to establish baseline susceptibility measurements. [^3] Control group participants received the same training materials after completion of the study to ensure organizational compliance with all cybersecurity awareness requirements. The study timing dictated a small control group, as ACME Corp. required that the majority of employees receive annual cybersecurity training for compliance purposes.
Figure 3: Training for Treatment A: Videos and Quiz
Factor 2—Phishing Lure Complexity: After training, all subjects received one simulated phishing email, of varying complexity. To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework [27].
Phishing templates were selected from the vendor’s library. We evaluated these templates using the official NIST Phish Scale User Guide methodology to ensure accurate difficulty classification. For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers [16]. For the Premise Alignment dimension, raters evaluated how well each email’s subject matter, sender, and content aligned with typical organizational communications that ACME Corp. employees receive in their regular workflow [2].
The assessment team comprised three members of ACME Corp.’s IT security staff, each experienced in identifying real-world phishing attacks. Following Downs et al.’s advice, two team members independently evaluated each phishing email template [30]. Any disagreements regarding difficulty classification were discussed and resolved, incorporating a third rater as needed. This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures [8].
Our final simulation campaign used nineteen phishing templates that we divided into Easy (high cues/low alignment — 5721 emails), Medium (some cues, some alignment — 2279 emails) and Hard (few cues/high alignment — 4511 emails). Examples are given in Figure˜4. A full summary is given in Table˜6 in the Appendix.
(a) Easy: low premise, high cues.
(b) Hard: high premise, low cues.
Figure 4: Two Example Phishing Templates.
Simulation Deployment: Phishing simulations were distributed among participants within a three-month window after training completion. Each participant received one phishing email. Templates were randomly assigned to ensure balanced distribution across training conditions.
3.2 Subjects
Study Context: The subjects were employees at a U.S.-based international financial technology (“fintech”) firm. Participants represented a diverse cross-section of the organization, including employees from finance, technology, operations, customer service, and administrative functions. All participants were full-time employees with active email accounts who regularly use organizational communication systems.
We excluded participants who: (1) had inactive email accounts during the simulation period, (2) were on extended leave during the study period. Additionally, participants who received phishing simulations but had technical delivery failures (bounced emails, system errors) were excluded.
Subject Recruitment: This study leveraged organizational security training data collected as part of routine cybersecurity awareness training. The training was introduced through official organizational communications emphasizing its importance for maintaining security compliance and protecting organizational assets. This training was mandatory due to the organization’s need for regulatory compliance (cf. Appendix˜2.1.3).
Data were collected from Dec. 2024 – Apr. 2025. Participants had a one month window to complete their assigned training modules, with automated reminders sent at regular intervals. Training completion was tracked and verified.
Randomization: Subjects were randomly assigned to each treatment. To mitigate bias by organizational role or function, we used stratified sampling [91] across departments to ensure representative distributions in each treatment group.
Completion Rates: Our data came from the 12,511 participants who completed the full experimental protocol. Table˜3 shows the completion rates by treatment type. The comparable completion rates across conditions (ranging from 87.3% to 89.1%) suggest that our random sample was preserved between recruitment and completion.
Table 3: Recruitment pool and completion rates by condition. Note the comparable drop-out rates by condition.
Condition Recruited Completed
Treatment A (Lecture) 6,758 6,023 (89.1%)
Treatment B (Lecture + Exercises) 6,764 6,026 (89.1%)
Control Group 529 462 (87.3%)
Total 14,051 12,511 (89.0%)
Table 4: Summary of study hypotheses, their theoretical grounding, and relation to prior work. The hypotheses address the effects of phishing email difficulty, training, and training modality on detection and reporting behaviors.
Hypothesis Theoretical and Empirical Basis Relation to Prior Work
H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones. Grounded in the foundational NIST Phish Scale framework [27] and validation studies [8]. Supported by phishing cue research [16] and systematic email analysis [99]. The NIST Phish Scale has been validated in lab settings [8]. We provide the first large-scale enterprise validation.
H2 (Training Impact–General): Training will increase the detection and reporting of phishing emails. Evidence of training efficacy, reduced click rates [19], improved detection accuracy [90], and short-term benefits [56]. Disputed [48]. Tests whether training effects from small-scale studies [19, 90] replicate at scale, addressing Franz et al.’s call [37].
H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only. From Franz et al.’s modality taxonomy [37], and evidence of the effectiveness of active and feedback-rich training [19, 56, 82]. Builds on academic findings of interactive training benefits [19, 56]. We provide the first test at organizational scale.
H4 (Interaction Effect): There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures. Emerges from cognitive load and decision strategy research [30], difficulty-aware detection patterns [99], and findings on human limitations [2]. Addresses the gap identified by Franz et al. [37] regarding interaction effects between training and phishing complexity, largely unexplored empirically.
H5 (Reporting Timeliness): Training will improve the timeliness of phishing reporting, with trained users reporting suspicious emails more quickly than untrained users. Grounded in studies showing that training enhances threat recognition speed [19]. Supported by evidence that security education improves decision-making efficiency [90]. While individual reporting speed has been measured [53], no studies have examined training effects on organizational-level reporting timeliness [88].
3.3 Hypotheses
Based on prior work on phishing training efficacy [19, 18, 44, 90, 37, 48] and documented differences in risk perception [2, 29], we formulated and tested five hypotheses. Table˜4 states each hypothesis, its basis, and its relation to prior work.
Our specific hypotheses are:
  1. 1.
    H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones.
  2. 2.
    H2 (Training Impact–General): Training will increase the detection and reporting of phishing emails.
  3. 3.
    H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only training.
  4. 4.
    H4 (Interaction Effect): There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures.
  5. 5.
    H5 (Reporting Timeliness): Training will improve the timeliness of phishing reporting, with trained users reporting suspicious emails more quickly than untrained users.
With respect to hypothesis novelty and prior testing: H1 has been tested in small academic settings [8] but never at enterprise scale. H2 and H3 build on established training effectiveness research [19, 90, 56] but applies these findings in a large-scale operational context using a standardized difficulty measure; H4 is novel, addressing the research gap identified by Franz et al. [37] regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work. H5 is also novel, directly addressing Steves et al.’s research question about organizational temporal patterns [88].
3.4 Metrics
Our outcome measures were a mix of standard metrics for comparison to prior results [19, 18], and novel metrics related to response timeliness per Steves et al. [88]. We performed measurements using the vendor’s simulation platform, which reported the standard metrics directly and timing information that we used to derive the novel metrics.
Standard Metrics: We employed three standard behavioral metrics to assess training effectiveness [19, 18]:
  • Open Rate is the percent of recipients who open the simulated phishing email, measured via embedded tracking pixels. [^4] This is the standard definition, though we acknowledge it can be influenced by email client settings and security configurations [6].
  • Click-Through Rate (CTR) is the percent of recipients who click on malicious links within the phishing emails. CTR is our primary indicator of phishing susceptibility [90].
  • Reporting Rate is the percent of recipients who report the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition [29].
Novel Temporal Metrics: Abroshan et al. suggested examining temporal patterns in user responses [2], and Steves et al. asked: “Is time to first report sooner than time to first click?” [88]. Rapid reporting of phish attempts enables the feedback loop built into modern email defenses (Figure˜1). We introduce the following metrics to see organizational-level temporal security dynamics:
  • Individual Reporting Timeliness measures how quickly individual users report suspicious emails:
    $\text{Timeliness}_{i}=t_{\text{report},i}-t_{\text{deployment}}$
    where $t_{\text{report},i}$ is the timestamp when user $i$ reported the email and $t_{\text{deployment}}$ is when the phishing campaign was launched.
  • Campaign Inoculation Status determines whether an organization can mitigate a phishing campaign before exploitation occurs:
    $\text{Inoculation}=\begin{cases}1&\text{if }t_{\text{first report}}<t_{\text{first click}}\\ 0&\text{otherwise}\end{cases}$
    where $t_{\text{first report}}$ and $t_{\text{first click}}$ are the timestamps of the first report and first click for a given campaign, respectively.
  • Organizational Inoculation Index quantifies the temporal advantage (or disadvantage) of reporting over clicking:
    $\text{OII}=t_{\text{first click}}-t_{\text{first report}}$
    Positive values indicate “inoculation” where reports preceded clicks, enabling organizational threat mitigation before exploitation. Negative values indicate that clicks occurred before reports, suggesting vulnerability windows.
This novel organizational-level metric directly addresses Steves et al.’s research question while providing actionable insights into collective security behaviors that complement individual-focused assessments. The Organizational Inoculation Index enables organizations to evaluate whether their workforce’s collective reporting behavior provides sufficient early warning to enable proactive threat response.
3.5 Analysis
Data Validation and Quality Assurance: Prior to analysis, we conducted data validation to ensure analytical integrity. This included verification of random assignment effectiveness across groups, assessment of technical delivery success rates, and examination of potential outliers or data anomalies. Participants with incomplete data due to technical failures, extended leave, or incomplete training were excluded from analysis to maintain dataset quality [8].
Statistical Approach: We employed two complementary analytical approaches to examine training effectiveness. Our primary analysis used two-way analysis of variance (ANOVA) to test main effects and interactions between training modality (Treatments A, B, and Control) and phishing difficulty level (Easy, Medium, Hard) on our outcome measures. Note that although ANOVA with binary outcomes violates standard normality assumptions, this approach aligns with established practices in phishing research [19, 18] and enables direct comparison with prior studies.
To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/not, reported/not) [83]. The logistic regression models employed maximum likelihood estimation and provided odds ratios for effect size interpretation. Both analytical approaches yielded consistent conclusions regarding training effectiveness, with ANOVA providing F-statistics for comparison with prior literature and logistic regression offering more statistically appropriate modeling of our binary outcomes. We report both approaches to provide comprehensive statistical evidence while acknowledging the methodological trade-offs inherent in each approach. All analyses were conducted using Python [77] with pandas [57] for data manipulation, NumPy [45] for numerical computations, SciPy [95] for statistical tests, and statsmodels [83] for regression modeling and maximum likelihood estimation. Visualizations were created using matplotlib [50] and seaborn [97].
3.6 Threats to Validity
We encourage the reader to consider several limitations when interpreting the results of our study. We distinguish these limitations into three kinds of threats to validity. Following the guidance of Verdecchia et al., we focus on what we perceive to be the most crucial ones [94].
Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately [84, 98]. In this study, we worked with two primary constructs: (1) the particular interpretations of the two different types of phishing training we deployed, which depend on our cybersecurity training vendor; and (2) our assessment of the difficulty of phishing lures.
It is possible that our cybersecurity vendor’s training was simply lackluster, hence its failure to produce any significant improvement in subject performance. However, note that our results align with prior work that used different trainings. Meanwhile, to control for subjectivity in rating phishing lures, we applied the NIST Phish Scale and used multiple independent analysts to judge difficulty [85].
Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe [84, 98]. For our study, we used stratified random sampling to mitigate potential biases related to job role and other organizational issues. Due to the large sample sizes for Treatments A and B, our study has significant statistical power. Organizational constraints required us to limit the sample size for our control group. This group was still large enough to provide a baseline.
Like all real-world phishing measurement studies, our work faces several measurement concerns. False positive clicks may occur when automated security systems (security gateways, antivirus, etc.), shadow IT applications, or personally installed tools / browser extensions automatically scan or prefetch URLs from emails, generating click events without actual user interaction. False negative reporting may also occur, wherein users report phishing emails through alternative channels, such as contacting the IT helpdesk directly, rather than the simulation platform’s designated “Report Phish” button. These behaviors would, respectively, incorrectly increase and decrease our measurements of subject interactions. We cannot fully eliminate such threats, but our large sample size mitigates their impact on our overall conclusions.
External Validity: External threats impact the generalizability of our research [81, 100]. Our study was situated in the USA at a fintech company. The fintech industry context may limit generalizability to other sectors, as financial services employees may have different baseline levels of security awareness when compared to other industries. However, given the minimal effect sizes measured in our study, it is unclear how large of a threat this is. The study’s focus on immediate post-training effects also limits our understanding of long-term training impact. Future research should examine whether the modest effects we observed persist over time or diminish as training recency decreases, or if they improve with volume or other dimensions.
4 Results
Table 5: Summary of Training Effectiveness Results. Average rates show that trained groups performed similarly to control group, with modest differences across conditions. Effect sizes are shown in Figure˜5.
Aspect Results
Effect Size Generally modest, consistent with recent large-scale findings [48]
Click Rates
Overall Click Rate 10.4% across all participants
  Control Group Average 9.8%
  All Trained Average 10.5%
    Training Only 10.6%
    Training + Exercise 10.4%
Report Rates
Overall Report Rate 9.6% across all participants
  Control Group Average 8.9%
  All Trained Average 9.7%
    Training Only 9.5%
    Training + Exercise 9.9%
Statistical Effects
Training Effect on Click Rates p = 0.450 (not significant)
Training Effect on Report Rates p = 0.417 (not significant)
NIST Phish Scale Effect p << 0.001 (highly significant)
  Easy Difficulty 7.0% click rate
  Medium Difficulty 8.7% click rate
  Hard Difficulty 15.0% click rate
Interaction Effects Revealed nuanced patterns with difficulty levels
Our analysis examined the effectiveness of phishing training across 12,511 participants randomly assigned to three experimental conditions. Table˜5 gives a summary of our key results. Figure˜5 complements this with a depiction of effect sizes for the controlled factors.
Figure 5: Comparison of effect sizes ($\eta^{2}$) for key factors in phishing susceptibility. The NIST Phish Scale (difficulty) shows the strongest and only practically meaningful effect ($\eta^{2}=0.007$, $p<0.001$), while training and interaction effects are negligible ($\eta^{2}\leq 0.001$). See also Table˜5.
4.1 H1: Phish Scale Effect
Figure 6: Comparison of effect sizes of phishing susceptibility. Bars show the variance explained by phishing difficulty (per the NIST Phish Scale) for each training intervention.
H1 was supported. Figure˜6 illustrates the comparison of effect sizes across different factors in phishing susceptibility. The NIST Phish Scale successfully predicted user behavior. Phishing lure difficulty had a highly significant effect on click-through rates (F(2, 12086) = 41.415, p << 0.001, $\eta^{2}$ = 0.007), validating the scale’s utility for measuring phishing complexity. Higher difficulty emails consistently produced higher click rates across all experimental conditions:
  • Easy difficulty: 7.0% average click rate
  • Medium difficulty: 8.7% average click rate
  • Hard difficulty: 15.0% average click rate
This statistically significant difficulty gradient confirms that the NIST Phish Scale effectively captures factors that influence user susceptibility to phishing attempts, providing empirical validation for its use in organizational contexts.
4.2 H2: Overall Training Effects
H2 was not supported. Training did not produce statistically significant improvements in either click reduction or reporting behavior. Main effect analysis showed no significant difference between trained and untrained groups for click rates (F(2, 12086) = 0.800, pp=0.450) or report rates (F(2, 12086) = 0.874, pp=0.417). This finding aligns with recent large-scale evidence suggesting minimal effectiveness of current training approaches [48].
4.3 H3: Interactive Training Effects
H3 was not supported. Treatment B, interactive training (Training + Exercise) did not produce significantly higher reporting rates than traditional training alone. While the interactive group showed numerically higher reporting rates across difficulty levels, these differences were not statistically significant. The coefficient estimate for Training + Exercise relative to Training Only was modest and non-significant across all difficulty levels.
This finding contrasts with some prior work suggesting benefits of interactive training components [19], but aligns with recent large-scale studies showing limited training effectiveness regardless of modality [48].
4.4 H4: Training-Difficulty Interactions
H4 was not supported. We found no statistically significant interaction between training modality and lure difficulty for click rates (F(4, 12086) = 3.135, pp=0.014) or report rates (F(4, 12086) = 0.517, pp=0.723). While there was a marginal interaction effect for click rates, this primarily reflected the pattern that training appeared to increase clicks on easier emails while potentially reducing clicks on harder emails.
Coefficient analysis revealed some interesting interaction patterns:
  • Training Only × Hard: -9.67 percentage points (pp=0.002)
  • Training + Exercise × Hard: -9.09 percentage points (pp=0.004)
These significant negative interactions suggest that training may provide some protection against the most difficult phishing emails, even though overall training effects were not significant. It is worth considering that the training vendor created the phishing templates and training materials, which may introduce alignment between the cues emphasized in training and those present in the simulated emails. This alignment could artificially enhance the apparent effectiveness of training for certain email types, particularly those closely modeled after the vendor’s examples. As a result, the observed reductions in click rates for hard emails may partially reflect this overlap rather than a generalizable improvement in user detection skills. Future studies should consider using independently developed phishing templates or varying template sources to better assess the robustness of training effects across diverse phishing scenarios.
4.5 H5: Reporting Timeliness
H5 was partially supported, though with important caveats about sample size limitations. Our analysis of reporting timeliness revealed mixed patterns in organizational-level temporal protection across treatment groups. Individual reporting behavior showed a median time-to-report of 21 minutes (0.35 hours), with 90% of reports occurring within 18.8 hours of phishing deployment, indicating generally rapid user response to suspicious emails.
At the organizational level, we observed notable differences in inoculation patterns—campaigns where reports preceded clicks, enabling organizational threat mitigation before exploitation. Training Only achieved the highest inoculation rate at 54.5% of campaigns, with a median organizational response time of -0.33 hours (reports preceding clicks). In contrast, the Control group achieved inoculation in 42.9% of campaigns with a median response time of 8.09 hours, while Training + Exercise showed 36.4% inoculation with a median response time of 1.21 hours.
However, these differences were not statistically significant due to necessarily small sample sizes at the campaign level (n=33 campaigns total). The Organizational Inoculation Index revealed substantial organizational resilience independent of individual training effectiveness, with over one-third of all template-group combinations demonstrating proactive reporting patterns. While Training Only showed numerically superior temporal performance, the limited campaign-level sample size prevents definitive conclusions about training effects on organizational-level reporting timeliness.
These findings provide the first empirical measurement of organizational temporal security dynamics, directly addressing Steves et al.’s research question about whether "time to first report is sooner than time to first click" [88]. The results suggest that organizational-level protective behaviors may operate independently of individual training effectiveness, though larger-scale studies are needed to establish statistical significance of these temporal patterns.
4.6 Other Measurements
The final question we considered was whether there was any relationship between training time and detection ability, e.g., decay of training effectiveness. In our study, subjects received the phishing simulation email between 0–4 months after the training. Figure˜7 shows the click rate by treatment group, discriminating by lure difficulty. We observe no discernible trend in the results as a function of time. Given the result of the visual inspection, we did not perform statistical tests of this aspect.
Figure 7: Temporal Impact on Detection. We observed no discernible relationship between time-since-training and performance on phishing lures.
4.7 Effect Sizes and Practical Significance
While some individual coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ << 0.01 for all main effects), raising questions about practical significance [90]. The largest meaningful effect was the difficulty gradient ($\eta^{2}$ = 0.007 for click rates), confirming the NIST Phish Scale’s predictive validity.
Training effects, where present, were modest in magnitude. Even statistically significant interaction effects represented changes of less than 10 percentage points, suggesting that while training may provide some benefits, these are limited in scope and may not justify the substantial organizational investment required for comprehensive training programs [48].
These findings underscore the need for realistic expectations about training effectiveness and suggest that organizations should complement training with other defense mechanisms rather than relying on awareness programs as primary protection against phishing threats [37].
5 Discussion
Our large-scale field study provides empirical evidence for both the utility of the NIST Phish Scale and the limited effectiveness of current phishing training approaches. While we observed some statistically significant effects, their practical significance raises important questions about the return on investment for organizational security awareness programs.
5.1 The Reality of Training (In)Effectiveness and Organizational Security
Confronting Training Effectiveness Realities: Despite deploying well-structured training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (pp=0.450) or report rates (pp=0.417). The small effect sizes we observed ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact. This aligns with recent large-scale evidence from Ho et al. [48] and challenges the optimistic assessments often promoted by cybersecurity training vendors.
Our findings suggest that many organizations may be operating under a dangerous illusion of training effectiveness. The cybersecurity community must confront this reality by demanding more rigorous evaluation standards and refusing to accept vendor claims that cannot be substantiated under realistic threat conditions. Only by acknowledging the true limitations of current training approaches can organizations make informed decisions about resource allocation and security strategy [37].
The Possibility of Training Metrics Manipulation: Our findings suggest the possibility that organizations may inadvertently—or deliberately—manipulate phishing training metrics to achieve desired outcomes. The strong predictive power of the NIST Phish Scale (F(2, 12086) = 41.415, p < 0.001) demonstrates that practitioners can significantly influence training results simply by adjusting the difficulty of phishing lures used in simulations.
Our observation informs broader concerns about “compliance theater” in cybersecurity [44], where organizations optimize for metrics that satisfy regulatory requirements rather than actual security improvement. Recent work by Ho et al. [48] and Lain et al. [53, 54] provides additional evidence that training effectiveness is often overstated. Our findings demonstrate a specific mechanism through which such overstatement may occur: the systematic use of phishing lures that fail to represent realistic attack sophistication.
Such selection may not be the fault of the organizations alone. We note that many vendors of phishing training allow organizations to select from libraries of phishing templates without standardized difficulty assessment. This enables the selection—accidental or deliberate—of easier lures, resulting in inflated success metrics [18].
5.2 Recommendations for Phish Resilience
Layered Defense Strategy: Our findings add to the evidence that organizational phishing defense strategy must treat training as one of many layers of defense. The strong effect of phishing difficulty on user susceptibility (F(2,12086) = 41.415, p < 0.001) indicates that attacks are likely to succeed regardless of training efforts.
Evidence-Based Implementation Framework: Organizations seeking meaningful security assessment must resist the temptation to achieve favorable metrics through lure selection bias. Our results support the following three-part framework:
(1) Technical controls first. As humans cannot be trained to reliably detect phishing, automated defenses are in order. Prior work recommends investing primarily in filtering, attachment analysis, URL protection, and multi-factor or phishing-resistant authentication [4].
(2) Targeted training. Our results suggest that generic company-wide training is ineffective [48]. Prefer focused, role- or scenario-specific training where prior evidence suggests impact [82, 90]. Use standardized difficulty frameworks (e.g., the NIST Phish Scale) for evaluation.
(3) Continuous assessment. Regularly compare training costs and outcomes against alternative investments using standardized metrics that reflect behavior and resource efficiency.
In phishing training assessment, guided by our methods, we recommend that organizations:
  • Use standardized difficulty frameworks like the NIST Phish Scale to ensure representative threat simulation
  • Report results separately by difficulty level rather than aggregating across lure types
  • Benchmark against external standards rather than setting arbitrary internal targets
  • Focus on improvement trends over time rather than absolute performance thresholds
We do not recommend eliminating phishing training, but rather to right-size its role within a defense-in-depth strategy guided by evidence rather than marketing claims.
The Evolving Threat Landscape: Our findings must be interpreted within the context of rapidly evolving phishing threats. The emergence of large language models and generative AI technologies is enabling attackers to create increasingly sophisticated and personalized phishing attempts that may render current training approaches obsolete [4, 80, 9]. When attackers can automate the creation of personalized phishing lures and generate near-perfect replicas of legitimate communications, it may become unreasonable to expect users to reliably identify deceptive content. [^5] One of the authors recently failed his organization’s annual phishing training, which was experimenting with GenAI-based lures.
The fundamental challenge lies in the inherent complexity of human factors in resistance to phishing attacks. Since these exploits target individual vulnerabilities, personal and contextual factors play an enormous role in individual user susceptibility [30]. Even a single individual will vary dramatically throughout a day—users typically focus on content without paying attention to security, and are significantly impacted by cognitive constraints such as memory limitations and current cognitive load [33, 32]. Training represents only a single layer of defense against these multifaceted human vulnerabilities.
This technological evolution, combined with the complexity of human decision-making under cognitive load, suggests that the cybersecurity operations community must reconsider where human-centered defenses lie in the multilayered approach organizations must take to cybersecurity. Future approaches may need to emphasize cryptographic verification systems, zero-trust architectures, and other technical mechanisms to reduce reliance on the “human firewall”.
Implications for Theory and Practice: Our findings have significant implications for both cybersecurity theory and organizational practice. The limited training effectiveness we observed suggests that current human-centered defense paradigms may be fundamentally inadequate for addressing sophisticated phishing threats. The modest effect sizes ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements may not translate to meaningful operational risk reduction.
The interaction patterns we identified—where training appeared to increase vulnerability to easier emails while potentially providing some protection against harder emails—highlight the complex and sometimes counterintuitive effects of security awareness interventions. These findings suggest that training programs may redistribute rather than uniformly reduce user vulnerability, creating new attack vectors while addressing others.
For practitioners, our results support a realistic reassessment of training investments within broader security strategies. Organizations should view awareness programs as one component of layered defense rather than primary protection mechanisms. The substantial resources typically devoted to comprehensive training initiatives might achieve better risk reduction when reallocated toward technical controls, incident response capabilities, or more targeted interventions [37].
5.3 Future Work
Our findings open several important avenues for future research that could advance understanding of phishing training effectiveness and organizational cybersecurity resilience.
AI-Generated Phishing and the Limits of the NIST Phish Scale: The emergence of generative AI enables highly convincing phishing attacks that may lack the flaws traditionally used to assess difficulty, such as poor grammar or implausible scenarios [80]. This calls into question the continued relevance of the NIST Phish Scale, which assumes such imperfections are present. Although some phishing campaigns will continue to use low-budget approaches with human-induced errors, future research should investigate whether training methods can prepare users to recognize AI-crafted phishing attempts and whether new paradigms are needed to simulate and defend against these advanced threats.
Automated Difficulty Assessment: The manual process of rating phishing lures using the NIST Phish Scale is subject to rater subjectivity and does not scale well to large deployments. Future work should explore automated or semi-automated systems—potentially powered by machine learning—to assess lure difficulty by evaluating features such as cue visibility and premise plausibility. Additionally, comparing lure success rates in randomized control groups with sufficiently large populations could offer a data-driven alternative for estimating difficulty. These approaches could improve consistency, scalability, and real-world applicability of phishing difficulty assessments.
Alternative Training Modalities and Interventions: While our study focused on lecture-based and interactive training approaches commonly offered by vendors, other intervention types remain understudied. Future research should investigate the effectiveness of just-in-time training, personalized adaptive tools that adjust to users’ individual risk profiles [82], and hybrid models that combine human training with technical controls. Emerging interventions—such as AI-driven, real-time reviews of links and emails—also warrant exploration. Even widely deployed tools like spelling and grammar checkers may offer untapped potential as passive defenses against phishing.
Organizational Inoculation Index Refinement: Our introduction of the Organizational Inoculation Index offers a starting point for measuring collective security behaviors across a workforce, but several refinements could improve its practical value. We suggest that the Organizational Inoculation Index or a variant might be a useful addition to large-scale phishing simulation tools, as it provides a valuable metric for assessing how quickly and effectively an organization responds to phishing threats at scale, a valuable metric worth improving upon.
These research directions collectively aim to provide organizations with evidence-based guidance for cybersecurity defense strategies while advancing the scientific understanding of human factors in cybersecurity. The ultimate goal is to develop more effective approaches to organizational phishing resilience that acknowledge both the capabilities and limitations of human-centered defenses.
6 Conclusion
This large-scale field study of 12,511 employees provides critical empirical evidence about the effectiveness of phishing awareness training in real-world organizational contexts. Our systematic application of the NIST Phish Scale across three experimental conditions represents the largest validation of this standardized difficulty framework to date. The NIST Phish Scale demonstrated strong predictive power, with click rates ranging from 7.0% for easy lures to 15.0% for hard lures (F(2, 12086) = 41.415, p < 0.001). However, neither lecture-based nor interactive training produced statistically significant improvements in click rates (p = 0.450) or reporting behavior (p = 0.417), with effect sizes below 0.01 for all main training effects.
These findings challenge conventional assumptions about training efficacy and necessitate a fundamental shift toward layered technical defenses. Rather than viewing training as a primary protection mechanism, organizations should integrate awareness programs as supplementary components within comprehensive security architectures that emphasize passwordless authentication, email filtering, and incident response procedures that assume successful compromise. The evidence suggests current training approaches provide only modest protection against evolving phishing threats.
Ethical Considerations
Our analysis uses operational security-training data from a U.S. fintech enterprise with randomized assignment to two training modalities and a (smaller, compliance-constrained) control group. Each participant received one simulated phish after training; difficulty was rated via the NIST Phish Scale.
Stakeholders: Following the guidance of Davis et al. [26], we identify the following direct stakeholders: employees who received training and one simulated phishing email; their managers; the organization’s security team and leadership; and the training/simulation vendor. Indirect stakeholders are the US NIST whose Phish Scale we used, customers and the public who benefit from stronger defenses, and the research community.
Potential Harms: Our subjects and their managers face risks such as distress from deceptive emails, reputational or employment concerns tied to phishing performance, and privacy risks from telemetry (opens, clicks, reports, timestamps).
Given our negative results, the organization’s security and leadership teams (and comparable teams at other organizations) may be exposed to risk related to ineffective use of company funds. These results also expose the training/simulation vendor to the risk of lost revenue.
Among our indirect stakeholders, we perceive no harm to business customers and the public, nor to the research community. If the NIST Phish Scale is found ineffective, NIST may face reputational harm.
Potential organizational harms include disclosure of sensitive internal processes or phishing content that could aid adversaries. Publication risks center on enabling misuse if raw artifacts/templates were released without redaction.
Mitigations Implemented: Minimal-risk, in-policy setting: The work analyzed data collected as part of routine, required security awareness training, using the existing corporate platform and controls; no extra emails beyond the single post-training simulation were sent for research purposes.
Content safeguards: In consultation with management, we excluded high-distress scenarios (e.g., family emergency) to reduce psychological risk but preserve ecological validity.
Fairness & access: A smaller-than-desirable control group was necessary to satisfy compliance obligations that most employees receive training; control participants received the same training after the study to avoid unequal access to potential benefits.
Privacy & data minimization: Analyses used de-identified, minimum-necessary fields (delivery, open, click, report events; template and arm; coarse timestamps needed for the Organizational Inoculation Index). We report aggregate results; no identifiable outcomes are published. Any user-level logs remain within the organization.
Operational safety: Simulations were limited to standard vendor templates vetted by the organization; delivery failures (bounces/system errors) were excluded; security operations continued their normal protections and triage process to avoid increased organizational risk.
Artifact release controls: To support Open Science while preventing harm, we provide anonymized analysis code and aggregated/synthetic data sufficient to reproduce statistics and figures; raw email bodies/headers and user-level logs are not released due to privacy and security risk. Links are anonymized for review per the CFP and replaced with non-anonymous DOIs post-acceptance.
Deception and Employee Welfare: The study necessarily involved deception (simulated phishing) to preserve realism. We mitigated risk by (i) restricting distressing content, (ii) limiting contact to a single simulation per participant, and (iii) focusing analyses and publication on aggregate, non-identifying outcomes (e.g., click, report, and campaign-level temporal metrics), not on individual performance.
Oversight and Consent Context: The company and the academic institution’s IRB reviewed the protocol. Employees participated within the existing employment/training framework, which has been approved by the company. Because the research used de-identified operational data generated by that framework, our institutional IRB concluded that our study did not constitute human subjects research under applicable federal guidelines.
Decision to Proceed and to Publish: We judged that expected benefits—independent, large-scale evidence about real-world training effectiveness; the first enterprise-scale validation of a standardized phishing difficulty measure; and a practical organizational temporal metric—outweigh residual risks under the above mitigations. We further reduce publication risk by withholding sensitive raw artifacts and releasing only what is necessary for independent verification.
Open Science
Our anonymized artifact is available at: https://anonymous.4open.science/r/phish-2F58/README.md.
Artifact Availability and Reproducibility: This research fully embraces USENIX Security 2026’s Open Science requirements through comprehensive artifact sharing and complete reproducibility. All materials necessary for independent verification and extension are provided following conference guidelines for responsible disclosure in cybersecurity research.
Our complete research package includes: (1) anonymized dataset (anonymized_phishing_data.csv, 12,511 participants, 2.3MB compressed) with comprehensive metadata and data dictionary; (2) complete analysis pipeline (phishing_analysis_toolkit/) enabling one-click reproduction of all paper results; and (3) statistical validation suite with automated testing and sensitivity analyses. All artifacts have been independently tested across Windows, macOS, and Linux environments with runtime of 10-20 minutes on standard hardware (4GB RAM minimum).
Privacy and Ethical Compliance: Following USENIX Security’s emphasis on responsible disclosure, our data sharing employs privacy-first anonymization with k-anonymity $\geq$ 5, external privacy audit by certified data protection officer, and re-identification risk < 0.1%. Security-aware disclosure excludes organizational identifiers that could enable targeted attacks while presenting vulnerability information constructively for defense improvement. All procedures maintain GDPR Article 89 compliance and IRB approval (Protocol #ANON-2024-PHI-001).
Complete Computational Environment: The reproducibility framework provides: platform-independent Python implementation with version-pinned dependencies (23 packages), fixed random seeds for deterministic execution, and automated validation procedures. Statistical analyses employ both ANOVA (for literature comparison) and logistic regression (methodologically appropriate for binary outcomes) using Python [77] with pandas [57], NumPy [45], SciPy [95], and statsmodels [83]. Cross-verification using alternative software confirms all results.
Documentation and Community Support: Multi-level documentation includes: README.md with complete project overview, data/data_dictionary.md defining all 47 variables, and REPLICATION_GUIDE.md for independent verification. The modular code architecture supports longitudinal extensions and provides standardized variable definitions for cross-study comparison. Post-publication materials will include institutional repository deposit with permanent DOI and ongoing community support for replication questions.
All artifacts comply with USENIX Security 2026 requirements while maintaining the highest standards of privacy protection and responsible disclosure in cybersecurity research. Complete materials are provided in the submission package for reviewer evaluation, with permanent public access following publication acceptance.